# Drinfeld rational fractions for affine Kac–Moody quantum symmetric pairs

### v2

@@ 7.2. Coproduct formula for $\boldsymbol{\grave{\Theta}}(z)$
(7.21) $\equiv$ $\ z^{2}c_{1}^{-1}C\left(z^{-1}A_{-1}\otimes[K,\boldsymbol{x}^{-}(-q^{-2}z^{-1})]_{q^{-2}}-q^{-2}A_{0}\otimes[K^{-1},\boldsymbol{x}^{-}(-q^{-2}z^{-1})]_{q^{2}}\right)$
@@
(7.26) $\equiv$ $\ -q^{-2}z^{2}c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\left(z^{-1}[x^{-}_{1},\boldsymbol{\phi}(-q^{-2}z^{-1})]_{q^{-2}}+[x^{-}_{0},\boldsymbol{\phi}(-q^{-2}z^{-1})]_{q^{2}}\right)$
@@
(7.29) $\equiv$ $\ z^{2}c_{1}^{-1}C(1-q^{-2})\boldsymbol{A}_{+}(z)\otimes\left(z^{-1}\xi_{\mathbf{c},\mathbf{s}}(A_{-1})K+\xi_{\mathbf{c},\mathbf{s}}(A_{0})K^{-1}\right)\boldsymbol{\phi}(-q^{-2}z^{-1}).$
@@
(7.31) $\equiv$ $\ z^{2}c_{1}^{-1}C(1-q^{-2})\boldsymbol{A}_{+}(z)\otimes\left(z^{-1}\xi_{\mathbf{c},\mathbf{s}}(A_{-1})K+\xi_{\mathbf{c},\mathbf{s}}(A_{0})K^{-1}\right)\boldsymbol{\phi}(-q^{-2}z^{-1}).$
@@
(7.33) $\equiv$ $\ -q^{-2}z^{2}c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\big{(}z^{-1}[x^{-}_{1},\kappa\Gamma]_{q^{-2}}+[x^{-}_{0},\kappa\Gamma]_{q^{2}}\big{)}$
(7.34) $\equiv$ $\ q^{-4}zc_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(-q^{-2}z^{-1})[\kappa\widetilde{\Gamma},(1-q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)x^{-}_{1}]$
(7.35) $\equiv$ $\ q^{-4}zc_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(-q^{-2}z^{-1})[(1+q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)\kappa\widetilde{\Gamma},x^{-}_{1}]$
(7.36) $\equiv$ $\ q^{-1}z^{2}(q-q^{-1})^{2}c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(-q^{-2}z^{-1})\cdot$
(7.37) $\cdot\big{[}(1+q^{-2}C\operatorname{ad}_{\overline{h}_{1}}z)^{-1}(\xi_{\mathbf{c},\mathbf{s}}(A_{0})x^{+}_{-1}-q^{-2}C\xi_{\mathbf{c},\mathbf{s}}(A_{-1})x^{+}_{0}),x^{-}_{1}\big{]}.$
Let us calculate the term in the last line above. Observe that
(7.38) $(1+q^{-2}C\operatorname{ad}_{\overline{h}_{1}}z)^{-1}(\xi_{\mathbf{c},\mathbf{s}}(A_{0})x^{+}_{-1}\ -$ $\ q^{-2}C\xi_{\mathbf{c},\mathbf{s}}(A_{-1})x^{+}_{0})=$
(7.39) $=$ $\ -q^{-2}C(\xi_{\mathbf{c},\mathbf{s}}(A_{-1})+\xi_{\mathbf{c},\mathbf{s}}(A_{0})z)\boldsymbol{x}^{+}_{\geq 0}(-q^{-2}Cz)$
(7.40) $+\xi_{\mathbf{c},\mathbf{s}}(A_{0})x^{+}_{-1}.$
@@
(7.41) $(q-q^{-1})\big{[}(1+q^{-2}C\operatorname{ad}_{\overline{h}_{1}}z)^{-1}($ $\xi_{\mathbf{c},\mathbf{s}}(A_{0})x^{+}_{-1}-q^{-2}C\xi_{\mathbf{c},\mathbf{s}}(A_{-1})x^{+}_{0}),x^{-}_{1}\big{]}=$
(7.42) $=$ $\ z^{-1}(\xi_{\mathbf{c},\mathbf{s}}(A_{-1})+\xi_{\mathbf{c},\mathbf{s}}(A_{0})z)\boldsymbol{\psi}_{\geq 1}(-q^{-2}Cz)$
(7.43) $+\xi_{\mathbf{c},\mathbf{s}}(A_{0})(K-K^{-1})$
(7.44) $=$ $\ (\xi_{\mathbf{c},\mathbf{s}}(A_{-1})z^{-1}+\xi_{\mathbf{c},\mathbf{s}}(A_{0}))\boldsymbol{\psi}(-q^{-2}Cz)$
(7.45) $-\xi_{\mathbf{c},\mathbf{s}}(A_{-1})z^{-1}K-\xi_{\mathbf{c},\mathbf{s}}(A_{0})K^{-1}.$
@@
(7.47) $\equiv$ $\ q^{-1}z^{2}(q-q^{-1})c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(-q^{-2}z^{-1})\cdot$
(7.48) $\{(\xi_{\mathbf{c},\mathbf{s}}(A_{-1})z^{-1}+\xi_{\mathbf{c},\mathbf{s}}(A_{0}))\boldsymbol{\psi}(-q^{-2}Cz)$
@@
(7.51) $+z(1-q^{-2})c_{1}^{-1}C(A_{-1}+A_{0}z)\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(-q^{-2}z^{-1})\boldsymbol{\psi}(-Cq^{-2}z).$
@@ 8.1. Quantum affine algebras
$\boldsymbol{\phi^{\pm}}(z)=\sum_{k=0}^{\infty}\phi^{\pm}_{i,\pm k}z^{\pm k}=K_{i}^{\pm 1}\exp\left(\pm(q-q^{-1})\sum_{k=1}^{\infty}h_{i,\pm k}z^{\pm k}\right).$
@@
$U_{d_{i}=r,+}=\langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\operatorname{deg}^{\mathrm{Dr}}_{i}y=r,\ \forall j\neq i\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y\geq 0,\ \exists j\neq i\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y>0\rangle.$
@@ 8.3. Generalized $q$-Onsager algebras
[…]
@@
(8.37) $B_{i}\mapsto A_{i,0},\quad\mathfrak{c}_{i}\mapsto\mathfrak{c}_{i},\quad\mathfrak{C}\mapsto q^{2N+2}\mathfrak{c}_{\delta},$
@@ 8.6. Rank one subalgebras
Let $U_{i}^{\mathbb{K}}$ be the subalgebra of $U_{q}^{\mathbb{K}}L\mathfrak{g}$ generated by $U_{i}$, $\mathbb{K}_{i}^{\pm 1}$ and $(\mathbb{K}_{\delta}\mathbb{K}_{i}^{-1})^{\pm 1}$. The algebra isomorphism $\iota_{i}$ extends to an isomorphism $\iota_{i}\colon U_{q}^{\mathbb{K}}L\mathfrak{sl}_{2}\to U_{i}^{\mathbb{K}}$ sending $\mathbb{K}_{1}\mapsto K_{i}$ and $\mathbb{K}_{0}\mapsto\mathbb{K}_{\delta}\mathbb{K}_{i}^{-1}$.
@@ 9.1. Action on generators
Assume that ${\mathfrak{g}}$ is not of type $\mathsf{A}_{1}$. We begin by investigating the image of elements such as $\mathbf{T}_{i}(B_{j})$ inside $U_{q}^{\mathbb{K}}L\mathfrak{g}$.
@@
At the level of generators, the image of the QSP braid group action is given by:
$\eta(\mathbf{T}_{i}(B_{j}))=\left\{\begin{array}[]{l l}\mathbb{K}_{j}^{-1}(F_{j}-q^{-2}\mathbb{K}_{j}E_{j}K_{j}^{-1})&\mbox{if }i=j,\\ (F_{j}-q^{-2}\mathbb{K}_{j}E_{j}K_{j}^{-1})&\mbox{if }a_{ij}=0,\\ T_{i}(\eta(B_{j}))+(q-q^{-1})q^{-2}\mathbb{K}_{j}F_{i}E_{j}K_{j}^{-1}&\mbox{if}a_{ij}=-1.\end{array}\right.$
@@
The first two cases follow directly from the definition. If $a_{ij}=-1$ then
@@
(9.8) $F_{j}E_{i}K_{i}^{-1}-qE_{i}K_{i}^{-1}F_{j}=0,$
concluding the proof. ∎
@@ 9.2. Type $\mathsf{A}$ combinatorics
From now on assume that ${\mathfrak{g}}$ is of type $\mathsf{A}_{N}$ with $N\geq 2$. Then $\Lambda=\langle\pi\rangle$, where $\pi$ is the automorphism of order $N+1$ of the affine Dynkin diagram of type $\mathsf{A}_{N}^{(1)}$ sending $i\mapsto i+1\mod N+1$. [Bou02, Ch. VI §2.3, Theorem 6] implies that the image of the fundamental weights under the isomorphism (8.18) can be described explicitly in the following way. Introduce a total order on the simple reflections $s_{i}$ ($i\in\mathbb{I}$) by setting: $s_{i}\leq s_{j}$ if and only if $i\leq j$. This induces a lexicographic partial order on words in $s_{i}$. Let $u=s_{1}s_{2}\cdots s_{N}$. By a subword of $u$ we mean a contiguous sequence of simple reflections. Let $S_{k}$ denote the set of subwords of $u$ of length $k$. The restriction of the lexicographic order to $S_{k}$ is a total order. Then, for any $i\in\mathbb{I}_{0}$, the image of the fundamental weight $\omega_{i}$ is given by the formula
$\omega_{i}\mapsto\pi^{i}\prod_{w\in S_{i}}^{\shortrightarrow}w,$
where the product is taken in order from the greatest word (left) to the lowest word (right). For example, let $N=5$. Then:
@@ 9.3. Inductive arguments
Let us set $\tilde{E}_{j}=-q^{-2}\mathbb{K}_{j}E_{j}K_{j}^{-1}$. We will need the following lemma.
@@
Let $2\leq m\leq N$. Then the element $\eta(\mathbf{T}_{m}\cdots\mathbf{T}_{2}(B_{1}))$ can be expressed as a sum of monomials consisting of letters $\tilde{E}_{1},\cdots,\tilde{E}_{l},F_{l+1},\cdots,F_{m}$ (with each letter occurring exactly once), where $0\leq l\leq m$. In particular, these monomials have degree $\sum_{j=1}^{l}\alpha_{j}-\sum_{j=l+1}^{m}\alpha_{j}$.
@@
We prove the statement by induction on $m$. The case $m=2$ reduces to Lemma 9.1. Otherwise, let us write
$\mathbf{T}_{m-1}\cdots\mathbf{T}_{2}(B_{1})=P_{m-1}(B_{1},\cdots,B_{m-1})$
as a polynomial in the letters $B_{1},\cdots,B_{m-1}$. Then
@@
(9.10) $=$ $\ P_{m-1}(B_{1},\cdots,B_{m-1})B_{m}-qB_{m}P_{m-1}(B_{1},\cdots,B_{m-1}),$
since $B_{m}$ commutes with $B_{1},\cdots,B_{m-2}$. By induction, $\eta(P_{m-1}(B_{1},\cdots,B_{m-1}))$ can be written as a polynomial $R_{m-1}(\bullet)$ in $\tilde{E}_{i},F_{i}$, which is a sum of monomials satisfying the conditions in the statement of the lemma. Hence
(9.11) $\eta(\mathbf{T}_{m}\cdots\mathbf{T}_{2}(B_{1}))=$ $\ R_{m-1}(\bullet)\eta(B_{m})-q\eta(B_{m})R_{m-1}(\bullet)$
(9.12) $=$ $\ R_{m-1}(\bullet;[F_{m-1},\eta(B_{m})]_{q};[\tilde{E}_{m-1},\eta(B_{m})]_{q})$
(9.13) $=$ $\ R_{m-1}(\bullet;[F_{m-1},F_{m}]_{q};[\tilde{E}_{m-1},\tilde{E}_{m}+F_{m}]_{q}),$
where $R_{m-1}(\bullet;a;b)$ denotes the polynomial $R_{m-1}(\bullet)$ with each occurrence of $F_{m-1}$ replaced by $a$, and each occurrence of $\tilde{E}_{m-1}$ replaced by $b$. The final equality follows from calculations similar to (9.5)–(9.8). Our calculation implies that the monomials in $\eta(\mathbf{T}_{m}\cdots\mathbf{T}_{2}(B_{1}))$ are obtained from those of $\eta(\mathbf{T}_{m-1}\cdots\mathbf{T}_{2}(B_{1}))$ by adding the letter $F_{m}$ if $F_{m-1}$ is already present, and letter $F_{m}$ or $\tilde{E}_{m}$ if $\tilde{E}_{m-1}$ is already present. Hence they have the required shape, which completes the proof. ∎
@@
Let $2\leq m\leq N$. Then the element $\eta(\mathbf{T}_{1}\cdots\mathbf{T}_{m-1}(B_{m}))$ can be expressed as a sum of monomials consisting of letters $\tilde{E}_{m},\cdots,\tilde{E}_{l},F_{l-1},\cdots,F_{1}$ (with each letter occurring exactly once), where $11\leq l\leq m+$ 1. In particular, these monomials have degree $-\sum_{j=1}^{l-1}\alpha_{j}+\sum_{j=l}^{m}\alpha_{j}$.

### v3

@@ 7.2. Coproduct formula for $\boldsymbol{\grave{\Theta}}(z)$
(7.18) $\equiv$ $\ z^{2}c_{1}^{-1}C\left(z^{-1}A_{-1}\otimes[K,\boldsymbol{x}^{-}(z^{-1})]_{q^{-2}}-q^{-2}A_{0}\otimes[K^{-1},\boldsymbol{x}^{-}(z^{-1})]_{q^{2}}\right)$
@@
(7.23) $\equiv$ $\ z^{2}c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\left(z^{-1}[x^{-}_{1},\boldsymbol{\phi}(z^{-1})]_{q^{-2}}-q^{-2}[x^{-}_{0},\boldsymbol{\phi}(z^{-1})]_{q^{2}}\right)$
@@
(7.26) $\equiv$ $\ z^{2}c_{1}^{-1}C(1-q^{-2})\boldsymbol{A}_{+}(z)\otimes\left(z^{-1}\xi_{\mathbf{c},\mathbf{s}}(A_{-1})K+\xi_{\mathbf{c},\mathbf{s}}(A_{0})K^{-1}\right)\boldsymbol{\phi}(z^{-1}).$
@@
(7.28) $\equiv$ $\ z^{2}c_{1}^{-1}C(1-q^{-2})\boldsymbol{A}_{+}(z)\otimes\left(z^{-1}\xi_{\mathbf{c},\mathbf{s}}(A_{-1})K+\xi_{\mathbf{c},\mathbf{s}}(A_{0})K^{-1}\right)\boldsymbol{\phi}(z^{-1}).$
@@
(7.30) $\equiv$ $\ z^{2}c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\big{(}z^{-1}[x^{-}_{1},\kappa\Gamma]_{q^{-2}}-q^{-2}[x^{-}_{0},\kappa\Gamma]_{q^{2}}\big{)}$
(7.31) $\equiv$ $\ -q^{-2}zc_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})[\kappa\widetilde{\Gamma},(1+q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x^{-}_{1}]$
(7.32) $\equiv$ $\ -q^{-2}zc_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})[(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\kappa\widetilde{\Gamma},x^{-}_{1}]$
(7.33) $\equiv$ $\ q^{-1}z^{2}(q-q^{-1})^{2}c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})\cdot$
(7.34) $\cdot\big{[}(1-C\operatorname{ad}_{\overline{h}_{1}}z)^{-1}(\xi_{\mathbf{c},\mathbf{s}}(A_{0})x^{+}_{-1}+C\xi_{\mathbf{c},\mathbf{s}}(A_{-1})x^{+}_{0}),x^{-}_{1}\big{]},$
where Lemma 4.3 was used in the third equality. Let us calculate the term in the last line above. Observe that
(7.35) $(1-C\operatorname{ad}_{\overline{h}_{1}}z)^{-1}(\xi_{\mathbf{c},\mathbf{s}}(A_{0})x^{+}_{-1}\ +$ $\ C\xi_{\mathbf{c},\mathbf{s}}(A_{-1})x^{+}_{0})=$
(7.36) $=$ $\ C(\xi_{\mathbf{c},\mathbf{s}}(A_{-1})+\xi_{\mathbf{c},\mathbf{s}}(A_{0})z)\boldsymbol{x}^{+}_{\geq 0}(Cz)+\xi_{\mathbf{c},\mathbf{s}}(A_{0})x^{+}_{-1}.$
@@
(7.37) $(q-q^{-1})\big{[}$ $(1-C\operatorname{ad}_{\overline{h}_{1}}z)^{-1}(\xi_{\mathbf{c},\mathbf{s}}(A_{0})x^{+}_{-1}+C\xi_{\mathbf{c},\mathbf{s}}(A_{-1})x^{+}_{0}),x^{-}_{1}\big{]}=$
(7.38) $=$ $\ z^{-1}(\xi_{\mathbf{c},\mathbf{s}}(A_{-1})+\xi_{\mathbf{c},\mathbf{s}}(A_{0})z)\boldsymbol{\psi}_{\geq 1}(Cz)+\xi_{\mathbf{c},\mathbf{s}}(A_{0})(K-K^{-1})$
(7.39) $=$ $\ (\xi_{\mathbf{c},\mathbf{s}}(A_{-1})z^{-1}+\xi_{\mathbf{c},\mathbf{s}}(A_{0}))\boldsymbol{\psi}(Cz)-\xi_{\mathbf{c},\mathbf{s}}(A_{-1})z^{-1}K-\xi_{\mathbf{c},\mathbf{s}}(A_{0})K^{-1}.$
@@
(7.41) $\equiv$ $\ q^{-1}z^{2}(q-q^{-1})c_{1}^{-1}C\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})\cdot$
(7.42) $\{(\xi_{\mathbf{c},\mathbf{s}}(A_{-1})z^{-1}+\xi_{\mathbf{c},\mathbf{s}}(A_{0}))\boldsymbol{\psi}(Cz)$
@@
(7.45) $+z(1-q^{-2})c_{1}^{-1}C(A_{-1}+A_{0}z)\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})\boldsymbol{\psi}(Cz).$
@@ 8.1. Quantum affine algebras
$\boldsymbol{\phi}^{\pm}_{i}(z)=\sum_{k=0}^{\infty}\phi^{\pm}_{i,\pm k}z^{\pm k}=K_{i}^{\pm 1}\exp\left(\pm(q-q^{-1})\sum_{k=1}^{\infty}h_{i,\pm k}z^{\pm k}\right).$
@@
(8.14) $U_{d_{i}=r,+}=$ $\ \langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\operatorname{deg}^{\mathrm{Dr}}_{i}y=r,\ \forall j\neq i\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y\geq 0,\ \exists j\neq i\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y>0\rangle,$
(8.15) $U_{d_{i}\geq r,+}=$ $\ \langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\operatorname{deg}^{\mathrm{Dr}}_{i}y\geq r,\ \forall j\neq i\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y\geq 0,\ \exists j\neq i\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y>0\rangle,$
@@
(8.17) $U_{+}=$ $\ \langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\forall j\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y\geq 0,\ \exists j\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y>0\rangle,$
(8.18) $U_{-}=$ $\ \langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\forall j\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y\leq 0,\ \exists j\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y<0\rangle,$
@@ 8.3. Generalized $q$-Onsager algebras
(8.37) $[A_{i,r},A_{j,s}]$ $=0\quad\mbox{if}\quad a_{ij}=0,$
(8.38) $[A_{i,r},A_{j,s+1}]_{q^{-a_{ij}}}-q^{-a_{ij}}[A_{i,r+1},A_{j,s}]_{q^{a_{ij}}}$ $=0\quad\mbox{if}\quad i\neq j,$
@@
(8.43) $B_{i}\mapsto A_{i,0},\quad\mathfrak{c}_{i}\mapsto\mathfrak{c}_{i},\quad q^{2N+2}\mathfrak{c}_{\delta}\mapsto\mathfrak{C},$
@@ 8.6. Rank one subalgebras
Let $U_{i}^{\mathbb{K}}$ be the subalgebra of $U_{q}^{\mathbb{K}}L\mathfrak{g}$ generated by $U_{i}$, $\mathbb{K}_{i}^{\pm 1}$ and $(\mathbb{K}_{\delta}\mathbb{K}_{i}^{-1})^{\pm 1}$. The algebra isomorphism $\iota_{i}$ extends to an isomorphism $\iota_{i}\colon U_{q}^{\mathbb{K}}L\mathfrak{sl}_{2}\to U_{i}^{\mathbb{K}}$ sending $\mathbb{K}_{1}\mapsto K_{i}$ and $\mathbb{K}_{0}\mapsto\mathbb{K}_{\delta}\mathbb{K}_{i}^{-1}$. For future use, we note that it follows easily from the definitions that
(8.45) $\iota_{i}(x^{-}_{r})=x^{-}_{i,r}.$
@@ 9.1. Action on generators
Assume that ${\mathfrak{g}}$ is not of type $\mathsf{A}_{1}$. Set
$\tilde{E}_{i}=-q^{-2}\mathbb{K}_{i}E_{i}K_{i}^{-1}.$
We begin by investigating the image of elements such as $\mathbf{T}_{i}(B_{i\pm 1})$ inside $U_{q}^{\mathbb{K}}L\mathfrak{g}$.
@@
If $a_{ij}=-1$ then
$\eta(\mathbf{T}_{i}(B_{j}))=T_{i}(\eta(B_{j}))+(q-q^{-1})F_{i}\tilde{E}_{j}=T_{i}(\eta(B_{j}))+[\tilde{E}_{j},F_{i}]_{q}.$
@@
If $a_{ij}=-1$ then
@@
(9.8) $F_{j}E_{i}K_{i}^{-1}-qE_{i}K_{i}^{-1}F_{j}=0.$
Applying (8.26) concludes the proof. ∎
For future use, let us record (9.8) as
(9.9) $[F_{i},\tilde{E}_{i\pm 1}]=0.$
@@ 9.2. Type $\mathsf{A}$ combinatorics
From now on assume that ${\mathfrak{g}}$ is of type $\mathsf{A}_{N}$ with $N\geq 2$. Then $\Lambda=\langle\pi\rangle$, where $\pi$ is the automorphism of order $N+1$ of the affine Dynkin diagram of type $\mathsf{A}_{N}^{(1)}$ sending $i\mapsto i+1\mod N+1$. The highest root is $\theta=\alpha_{1}+\alpha_{2}+\cdots+\alpha_{n}$.
Let $w_{0}$ be the longest element of the finite Weyl group $W_{0}$ of ${\mathfrak{g}}$. Given $i\in\mathbb{I}_{0}$, let $w_{i}$ the longest element of the Weyl group corresponding to the Dynkin diagram of ${\mathfrak{g}}$ with the node labelled by $i$ removed. Given $1\leq k<l$, let $[k,l]=s_{k}s_{k+1}\cdots s_{l}$.
Proposition 9.2.
Let $\widetilde{W}$ be the extended affine Weyl group of type $\mathsf{A}_{N}^{(1)}$, for $N\geq 1$. Then
$\omega_{i}=\pi^{i}[N-i+1,N]\cdots[2,i+1][1,i].$
This is a reduced expression.
Proof.
It follows from [Bou02, Ch. VI §2.3, Theorem 6] that
$\omega_{i}=\pi^{i}w_{0}w_{i}.$
It is easy to check that $w_{0}w_{i}=\tau^{N+1-i}$, where $\tau$ is the inverse of the cyclic permutation $(12\cdots N+1)$, if we identify $W$ with the symmetric group $S_{N+1}$. It is well known that such permutations have an expression of the form $[N-i+1,N]\cdots[2,i+1][1,i]$. To show that this expression is reduced, it suffices to check that $\ell(\omega_{i})=i(N-i+1)$. By [IM65, Proposition 1.23], $\ell(\omega_{i})$ is equal to the number of positive roots which contain $\alpha_{i}$ in their expression in terms of the basis of simple roots. Such roots are precisely $\alpha_{k}+\cdots+\alpha_{l}$, for $1\leq k\leq i$ and $i\leq l\leq N$, completing the proof. ∎
For example, let $N=5$. Then:
@@ 9.3. Inductive arguments
Define non-commutative polynomials $P_{k}(y_{1},\cdots,y_{k})$ over ${\mathbb{C}}$ by induction in the following way:
(9.10) $P_{1}(y_{1})=$ $\ y_{1},\qquad$ $P_{k+1}(y_{1},\cdots,y_{k})=$ $\ [P_{k}(y_{1},\cdots,y_{k}),y_{k+1}]_{q}.$
Clearly, we have
(9.11) $P_{k}(y_{1},\cdots,y_{k})=$ $\ P_{k-l+1}(P_{l}(y_{1},\cdots,y_{l}),y_{l+1},\cdots,y_{k}).$
for any $11\leq l\leq k-$ 1. We will need the following lemma.
@@
Let $2\leq m\leq N$. Then $\mathbf{T}_{m}\cdots\mathbf{T}_{2}(B_{1})=P_{m}(B_{1},\cdots,B_{m})$. Moreover,
$\eta(\mathbf{T}_{m}\cdots\mathbf{T}_{2}(B_{1}))=\sum_{\begin{subarray}{c}\varepsilon_{k}\geq\varepsilon_{k+1},\\ k=1,...,m-1\end{subarray}}P_{m}(\tilde{e}_{1}^{\varepsilon_{1}},\cdots,\tilde{e}_{m}^{\varepsilon_{m}}),$
where each $\varepsilon_{l}\in\{+,-\}$, and $\tilde{e}_{i}^{+}=\tilde{E}_{i}$, $\tilde{e}_{i}^{-}=F_{i}$. In other words, the RHS consists of polynomials in letters $\tilde{E}_{1},\cdots,\tilde{E}_{l},F_{l+1},\cdots,F_{m}$ (with each letter occurring exactly once), for $0\leq l\leq m$. In particular, these polynomials have degree $\sum_{j=1}^{l}\alpha_{j}-\sum_{j=l+1}^{m}\alpha_{j}$.
@@
We prove the statement by induction on $m$. The case $m=2$ reduces to Lemma 9.1. Otherwise, by induction,
$\mathbf{T}_{m-1}\cdots\mathbf{T}_{2}(B_{1})=P_{m-1}(B_{1},\cdots,B_{m-1}).$
Then
@@
(9.13) $=$ $\ [P_{m-1}(B_{1},\cdots,B_{m-1}),B_{m}]_{q}$
(9.14) $=$ $\ P_{m}(B_{1},\cdots,B_{m}),$
since $B_{m}$ commutes with $B_{1},\cdots,B_{m-2}$. This proves the first statement of the lemma.
Moreover, by induction,
$\eta(P_{m-1}(B_{1},\cdots,B_{m-1}))=\sum_{\varepsilon_{k}\geq\varepsilon_{k+1}}P_{m-1}(\tilde{e}_{1}^{\varepsilon_{1}},\cdots,\tilde{e}_{{m-1}}^{\varepsilon_{m-1}}).$
Therefore, it suffices to show that
$P_{m}(\tilde{e}_{1}^{\varepsilon_{1}},\cdots,\tilde{e}_{{m-2}}^{\varepsilon_{m-2}},F_{m-1},\tilde{E}_{m})=0.$
But this is indeed the case since
$P_{m}(\tilde{e}_{1}^{\varepsilon_{1}},\cdots,\tilde{e}_{{m-2}}^{\varepsilon_{m-2}},F_{m-1},\tilde{E}_{m})=P_{m-1}(\tilde{e}_{1}^{\varepsilon_{1}},\cdots,\tilde{e}_{{m-2}}^{\varepsilon_{m-2}},[F_{m-1},\tilde{E}_{m}]_{q})$
and $[F_{m-1},\tilde{E}_{m}]_{q}=0$ by (9.9). ∎
@@
Let $2\leq m\leq N$. Then $\mathbf{T}_{1}\cdots\mathbf{T}_{m-1}(B_{m})=P_{m}(B_{m},\cdots,B_{1})$. Moreover,
$\eta(\mathbf{T}_{1}\cdots\mathbf{T}_{m-1}(B_{m}))=\sum_{\begin{subarray}{c}\varepsilon_{k}\leq\varepsilon_{k+1},\\ k=1,...,m-1\end{subarray}}P_{m}(\tilde{e}_{m}^{\varepsilon_{m}},\cdots,\tilde{e}_{1}^{\varepsilon_{1}}),$
In other words, the RHS can be expressed as a sum of polynomials consisting of letters $\tilde{E}_{m},\cdots,\tilde{E}_{l},F_{l-1},\cdots,F_{1}$, for $11\leq l\leq m+$ 1. In particular, these polynomials have degree $-\sum_{j=1}^{l-1}\alpha_{j}+\sum_{j=l}^{m}\alpha_{j}$.